# Kernelized Decoded Quantum Interferometry

### v2

@@ V Circuit cost and efficiency
V.1 Resource model and module costs
@@
Lemma V.1 (Synthesis of quadratic chirp kernels)
@@
V.2 End-to-end scaling and block-local advantage
Circuit resources for one k-DQI iteration.
@@ V.2 End-to-end scaling and block-local advantage
$\textnormal{two-qubit depth}\;=\;\tilde{O}\!\big(\ell n\big)\;+\;\tilde{O}\!\big(\text{depth}(K)\big)\;+\;\tilde{O}\!\big(\text{depth}(\mathcal{F})\big)\;+\;\tilde{O}\!\big(\text{depth}(Dec_{\mathrm{coh}})\big),$ (22)
$T\textnormal{-count}\;=\;\tilde{O}\!\big(\ell n\big)\;+\;\tilde{O}\!\big(T(K)\big)\;+\;\tilde{O}\!\big(T(\mathcal{F})\big)\;+\;\tilde{O}\!\big(T(Dec_{\mathrm{coh}})\big),$ (23)
@@ V.3 Cost–benefit frontier and scaling plots
Cost–benefit frontier.
Figure 6 summarizes the empirical cost–benefit trade-off for the OPI setting. The horizontal axis reports relative two-qubit depth, normalized so that standard DQI with identity kernel has cost 1, while the vertical axis shows the head-mass gain $\Sigma_{K}(\ell,\eta;d_{\star})-\Sigma_{I}(\ell,\eta;d_{\star})$ computed from the same polynomial-phase model as in Figure 3. The baseline point $(1,0)$ corresponds to unkernelized DQI. A cloud of mis-tuned global chirp kernels (gray circles) sits at essentially the same relative depth as the tuned chirp, but with negligible head-mass gain. In contrast, a tuned global chirp (black diamond) delivers a large gain at this higher cost, while block-local chirp kernels (colored stars) achieve moderate gains at near-baseline depth. Together with the monotone and noise-aware bounds in Sections II.2 and III, this figure highlights that the improvement of k-DQI over DQI is driven by structure-adaptive kernels—global or block-local—rather than by a mere increase in circuit depth.
@@
Depth scaling with problem size.
Figure 7 summarizes these depth models. On the log–log plot, k-DQI with global chirp/LCT kernels rapidly peels away from the baseline and approaches the $n^{2}$ reference line, reflecting the quadratic cost of a fully global kernel layer. By contrast, unkernelized DQI and k-DQI with block-local kernels of fixed width $b=O(1)$ remain parallel to the $n$ reference line. The inset makes this behaviour more transparent: for all $n$ in the plotted range, block-local kernels lie only slightly above the baseline curve, with an overhead of about a constant factor that does not grow with $n$. Together with the cost–benefit frontier in Figure 6, this shows that the most practical structure-adaptive gains of k-DQI come from block-local kernels, which preserve near-linear two-qubit depth while still delivering the head-mass improvements certified in Sections IV and III.
Figure 7: Asymptotic two-qubit depth per iteration for DQI and k-DQI on a log–log scale. The main panel plots the depth models from Section V as a function of problem size $n$: baseline DQI with $K=I$ (solid black), k-DQI with block-local kernels of width $b\in\{4,8\}$ (blue and orange), and k-DQI with global chirp/LCT kernels (black dashed). Gray curves show reference scalings proportional to $n$ and $n^{2}$. The global kernel quickly follows the quadratic reference, while the baseline and block-local curves track the linear reference. The inset zooms into the near-linear region, making it clear that block-local kernels incur only a mild constant-factor overhead over unkernelized DQI while preserving essentially linear scaling in $n$.
V.4 Accuracy and synthesis overhead
@@
Kernels compatible with our analysis (chirp/LCT and block-local variants) are implementable with polynomial resources and, in the block-local regime, preserve the near-linear depth scaling of the original DQI architecture. Thus the structure-adaptive improvements certified by Theorems IV.1 and IV.2 are not only information-theoretic, but also architecturally consistent with realistic circuit models.
VI Scalability of Kernel Parameter Search
@@
VI.1 Regime I: Explicit Algebraic Structure
For problems defined by known polynomial constraints, such as the Optimal Polynomial Interpolation (OPI) problem discussed in Section IV, the kernel parameters are analytically derivable. Given a target polynomial $\mathcal{P}$ of degree $\ell$, the optimal spectral shaping is achieved by matching the quadratic phase profile of the kernel to the curvature of $\mathcal{P}$. specifically, for a chirp kernel $K(\gamma)=\exp(-i\gamma\hat{x}^{2})$, the optimal rate satisfies $\gamma^{*}\propto\text{coeff}(\mathcal{P})$, which can be computed in $O(1)$ classical time. In this regime, the “search” cost is zero, and k-DQI provides a deterministic compilation advantage.
VI.2 Regime II: Variational k-DQI for Latent Structure
@@
Figure 8: Optimization landscape of the multi-parameter kernel Ansatz. (a) Surface topography of the noise-weighted head mass $\Sigma_{K}(\theta_{1},\theta_{2})$ as a function of chirp rate $\theta_{1}$ and mixing angle $\theta_{2}$. The simulation incorporates local depolarizing noise ($\eta=0.1$) and finite sampling shot noise ($N_{shots}=5000$), resulting in the rugged surface texture. Despite coherent sidelobes (ripples) and statistical fluctuations, the landscape exhibits a dominant, convex-like basin of attraction. (b) Contour projection overlaid with a variational gradient descent trajectory (black/white path). Starting from a random initialization with negligible overlap (white circle, $\Sigma_{K}<0.1$), the classical optimizer reliably navigates the gradient flow to the global optimum (red star, $\Sigma_{K}\approx 0.92$) within 40 iterations. (Inset) The learning curve showing monotonic improvement of the objective function. This empirically confirms the smoothness guarantees derived in Appendix C, demonstrating that k-DQI spectral concentration is learnable efficiently even in the absence of precise a priori structural knowledge.
Crucially, our data indicates that $\Sigma_{K}$ is robust to parameter misspecification: achieving 90% of the peak advantage only requires precision $\delta\theta\sim O(1/\sqrt{n})$, as analyzed in Appendix E.
VII Discussion and outlook
@@
From a complexity standpoint, Section V shows that our kernels admit faithful circuit models with poly resources, both in global and block-local forms. the same kernels that certify a lift in $\Sigma_{K}$ can, in principle, be realized coherently alongside reversible decoders such as Berlekamp–Massey. While our guarantees do not require coherent decoders—classical post-processing suffices—the architectural consistency makes the results actionable for future hardware studies.
@@
Acknowledgements.
We acknowledge the support of the National Nature Science Foundation of China (Grant No. 12174301), the Natural Science Basic Research Program of Shaanxi (Grant No. 2023-JC-JQ-01), the Open Fund of State Key Laboratory of Acoustics (Grant No. SKLA202312), the Shaanxi Province Postdoctoral Science Foundation (Grant No.2024BSHEDZZ021), the Aeronautical Science Fund of China(Grant No. 2024M075070001), and the Project of Institute of Medical Intelligence of the First Affiliated Hospital of Xi’an Jiaotong University.
Appendix A Proofs for Section II
We supply the technical ingredients used in II.1. Throughout, $g(x)=P\!\bigl(f(x)\bigr)$, $\alpha=\mathcal{F}K\,g$, $\mathcal{S}_{d}=\mathcal{S}_{d}(\alpha)$, and $\Sigma_{K}(\ell,\eta;d)$ is as in (4).
@@ A.1 Noise contraction and leakage
Lemma A.1 (Per‑mode contraction under local noise)
@@
proof A.1
Local depolarizing contracts off‑diagonal terms in the post‑interference density matrix by $(1-\eta)$ per qubit, which translates to a per‑mode attenuation on diagonal elements (Born probabilities) bounded below by $(1-\eta)^{\kappa(s)}$ for some weight proxy $\kappa(s)$; nonuniform loss further rescales by $\tau(s)$. Banded mixing $\mathcal{M}$ can be written as a unitary that redistributes probability mass among indices in a ww‑neighborhood; by unitarity and bounded mixing width, the loss of mass from $\mathcal{S}_{d}$ is at most a function $\Delta(w)$ that vanishes with $w$. Combining these gives the stated inequality (cf. [Bu2025Noise] for similar contractions at the level of interference fringes).
@@ A.2 Head–tail interference and decoder coupling
Lemma A.2 (Head–tail coherence bound)
@@
proof A.2
Write $\alpha=(\mathcal{F}K)g$ and decompose $g=g_{\mathrm{head}}+g_{\mathrm{tail}}$ along the preimage of $\mathcal{S}_{d}$ under $(\mathcal{F}K)^{\ast}$. The off‑support leakage is governed by $\|\Pi_{\mathcal{S}_{d}^{c}}(\mathcal{F}K)\Pi_{\mathcal{S}_{d}}\,g_{\mathrm{head}}\|_{2}$, which is $\leq\mu\|g\|_{2}$ by definition of $\mu$; squaring gives the probability bound.
Proposition A.1 (Decoder threshold coupling)
@@
A.3 Proof of Theorem II.1
proof A.3 (Proof of II.1)
Apply A.1 to lower‑bound the post‑channel head probability by $\Sigma_{K}(\ell,\eta;d)-\Delta(w)$, then subtract the head–tail leakage from A.2 if desired (this can be absorbed into $\Delta(w)$, or treated separately). By A.1 and the monotonicity of $\Phi$,
@@
Finally, the DQI reduction maps decoder success to an approximation ratio lower bound $A(\ell,\eta,K;d)$ (see [JordanNature2025]); since $\Phi$ is nondecreasing in its first argument and $\Delta(w)$ does not depend on $K$, the right‑hand side is a nondecreasing function of $\Sigma_{K}(\ell,\eta;d)$. If two kernels give strictly ordered head masses above threshold, the inequality is strict by the monotonicity of $\Phi$, establishing the claim.
@@ Appendix B Noise model and head-mass bounds
In this appendix we collect the technical bounds that relate the physical noise model $\mathcal{E}=\mathcal{L}\circ\mathcal{M}\circ\mathcal{Z}$ to the noise-weighted head-spectrum mass $\Sigma_{K}(\ell,\eta;d)$ defined in Section II.2. Throughout we use the notation $g(x)=P(f(x))$, $\alpha=\mathcal{F}K\,g$, and $\mathcal{S}_{d}(\alpha)$ for the indices of the $d$ largest $|\alpha_{s}|$.
@@ B.1 Standing assumptions
    The decoder $\mathrm{Dec}$ satisfies the monotone threshold property described in Section II.3: its success probability for achieving a target approximation ratio $\rho$ is a nondecreasing function of the noise-weighted head-spectrum mass $\Sigma_{K}(\ell,\eta;d)$.
@@ B.2 Per-mode contraction and leakage
Lemma B.1 (Per-mode contraction)
@@
with $\Delta\to 0$ as $W\to 0$. This is exactly the claimed inequality.
@@ B.3 Head–tail coherence and effective head probability
Lemma B.2 (Head–tail coherence)
@@
proof B.2
Write the matrix of $\mathcal{F}K$ in block form $M=\begin{pmatrix}A&B\\ C&D\end{pmatrix}$ with respect to the head/tail decomposition. The tail component of $\alpha$ is $Cg_{\mathrm{head}}+Dg_{\mathrm{tail}}$. By assumption, $\|C\|_{2\to 2}\leq\mu$. Using $\|Cg_{\mathrm{head}}\|_{2}\leq\mu\|g_{\mathrm{head}}\|_{2}$ and unitarity of $M$ (which implies $\|D\|\leq 1$), we obtain $\|\alpha_{\mathrm{tail}}\|_{2}\leq\mu\|g_{\mathrm{head}}\|_{2}+\|g_{\mathrm{tail}}\|_{2}\leq(\mu^{2}+1)\|g\|_{2}$ and hence the stated bound up to a constant factor. A sharper estimate can be obtained by orthogonal decomposition of the range of $C$ and absorbing constants into $\mu$.
@@
Proposition B.1 (Effective head probability)
@@
proof B.3
Lemma B.1 lower-bounds the head probability after noise but before taking into account coherent leakage; Lemma A.2 bounds the additional probability that can move from the head into the tail due to head–tail blocks of $\mathcal{F}K$. Subtracting the coherence penalty from the right-hand side of the inequality in Lemma B.1 gives the claim.
@@ B.4 Noise-aware lower bound
We can now restate the noise-aware lower bound used in Section III in a self-contained way.
Theorem B.1 (Noise-aware lower bound)
@@
proof B.4
By Proposition B.1, the post-channel head probability is at least the effective head mass $\Sigma_{K}(\ell,\eta;d)-\Delta(w)-\mu^{2}\|g\|_{2}^{2}$. The decoder monotonicity property (assumption (A3) and Section II.3) converts any increase in this effective head mass into a nondecreasing improvement of the decoder success probability. The standard DQI reduction from decoding success to objective value (e.g. [JordanNature2025]) then yields the lower bound $A(\ell,\eta,K;d)$ with a response curve $F$ that depends on the decoder family and problem ensemble (BM/KV/GS for RS-like codes [GuruswamiSudan1999, KoetterVardy2003], density evolution / EXIT for LDPC-like ensembles [RichardsonUrbanke2001, RichardsonShokrollahiUrbanke2001, Kschischang2001Factor]). Monotonicity of $F$ in its second argument gives the final statement.

### v3

@@ 3 Methods
3.1 Resource model and module costs
@@
Lemma 3.1 (Synthesis of quadratic chirp kernels).
@@
3.2 End-to-end scaling and block-local advantage
@@
2-qubit depth $=\tilde{O}\!\big(\ell n\big)+\tilde{O}\!\big(\text{depth}(K)\big)+\tilde{O}\!\big(\text{depth}(\mathcal{F})\big)+\tilde{O}\!\big(\text{depth}(Dec_{\mathrm{coh}})\big),$ (22a)
$T\textnormal{-count}$ $=\tilde{O}\!\big(\ell n\big)+\tilde{O}\!\big(T(K)\big)+\tilde{O}\!\big(T(\mathcal{F})\big)+\tilde{O}\!\big(T(Dec_{\mathrm{coh}})\big).$ (22b)
@@
Figure 7 summarizes the empirical cost–benefit trade-off for the OPI setting. The horizontal axis reports relative two-qubit depth, normalized so that standard DQI with identity kernel has cost 1, while the vertical axis shows the head-mass gain $\Sigma_{K}(\ell,\eta;d_{\star})-\Sigma_{I}(\ell,\eta;d_{\star})$ computed from the same polynomial-phase model as in Figure 3. The baseline point $(1,0)$ corresponds to unkernelized DQI. A cloud of mis-tuned global chirp kernels (gray circles) sits at essentially the same relative depth as the tuned chirp, but with negligible head-mass gain. In contrast, a tuned global chirp (black diamond) delivers a large gain at this higher cost, while block-local chirp kernels (colored stars) achieve moderate gains at near-baseline depth. Together with the monotone and noise-aware bounds in Sections 2.2 and 2.4, this figure highlights that the improvement of k-DQI over DQI is driven by structure-adaptive kernels—global or block-local—rather than by a mere increase in circuit depth.
@@
Figure 8 summarizes these depth models. On the log–log plot, k-DQI with global chirp/LCT kernels rapidly peels away from the baseline and approaches the $n^{2}$ reference line, reflecting the quadratic cost of a fully global kernel layer. By contrast, unkernelized DQI and k-DQI with block-local kernels of fixed width $b=O(1)$ remain parallel to the $n$ reference line. The inset makes this behaviour more transparent: for all $n$ in the plotted range, block-local kernels lie only slightly above the baseline curve, with an overhead of about a constant factor that does not grow with $n$. Together with the cost–benefit frontier in Figure 7, this shows that the most practical structure-adaptive gains of k-DQI come from block-local kernels, which preserve near-linear two-qubit depth while still delivering the head-mass improvements certified in Sections 2.5 and 2.4.
Figure 8: Asymptotic two-qubit depth per iteration for DQI and k-DQI on a log–log scale. The main panel plots the depth models from Section 3 as a function of problem size $n$: baseline DQI with $K=I$ (solid black), k-DQI with block-local kernels of width $b\in\{4,8\}$ (blue and orange), and k-DQI with global chirp/LCT kernels (black dashed). Gray curves show reference scalings proportional to $n$ and $n^{2}$. The global kernel quickly follows the quadratic reference, while the baseline and block-local curves track the linear reference. The inset zooms into the near-linear region, making it clear that block-local kernels incur only a mild constant-factor overhead over unkernelized DQI while preserving essentially linear scaling in $n$.
@@
Kernels compatible with our analysis (chirp/LCT and block-local variants) are implementable with polynomial resources and, in the block-local regime, preserve the near-linear depth scaling of the original DQI architecture. Thus the structure-adaptive improvements certified by Theorems 2.5 and 2.5 are not only information-theoretic, but also architecturally consistent with realistic circuit models.
3.3 Scalability of Kernel Parameter Search
@@
For problems defined by known polynomial constraints, such as the Optimal Polynomial Interpolation (OPI) problem discussed in Section 2.5, the kernel parameters are analytically derivable. Given a target polynomial $\mathcal{P}$ of degree $\ell$, the optimal spectral shaping is achieved by matching the quadratic phase profile of the kernel to the curvature of $\mathcal{P}$. specifically, for a chirp kernel $K(\gamma)=\exp(-i\gamma\hat{x}^{2})$, the optimal rate satisfies $\gamma^{*}\propto\text{coeff}(\mathcal{P})$, which can be computed in $O(1)$ classical time. In this regime, the “search” cost is zero, and k-DQI provides a deterministic compilation advantage.
@@
Figure 9: Landscape of a two-parameter kernel Ansatz in k-DQI. (a) Noise-weighted head-mass landscape $\Sigma_{K}(\theta_{1},\theta_{2})$ for a two-parameter kernel $K(\theta_{1},\theta_{2})$ composed of a quadratic-phase chirp (parameter $\theta_{1}$) followed by a nearest-neighbour mixing layer (parameter $\theta_{2}$). The landscape is evaluated for a small k-DQI instance under local depolarizing noise, and $\Sigma_{K}$ is estimated from finite-shot sampling, which produces the fine-scale roughness visible on top of an otherwise broad, single-basin structure. (b) Contour representation of the same landscape, overlaid with a gradient-ascent trajectory (black/white curve) from a random initialization (white marker) to a high-mass region (red marker) in the kernel space. The inset shows the value of $\Sigma_{K}$ along the trajectory, exhibiting a steady improvement of the objective up to small stochastic fluctuations. Together, these panels illustrate that in the low-dimensional kernel parameter space the effective objective for k-DQI is well behaved and amenable to simple gradient-based search even in the presence of noise and sampling effects.
4 Discussion
@@
From a complexity standpoint, Section 3 shows that our kernels admit faithful circuit models with poly resources, both in global and block-local forms. the same kernels that certify a lift in $\Sigma_{K}$ can, in principle, be realized coherently alongside reversible decoders such as Berlekamp–Massey. While our guarantees do not require coherent decoders—classical post-processing suffices—the architectural consistency makes the results actionable for future hardware studies.
@@
[…]
Acknowledgements
We acknowledge the support of the National Nature Science Foundation of China (Grant No. 12174301), the Natural Science Basic Research Program of Shaanxi (Grant No. 2023-JC-JQ-01), the Open Fund of the State Key Laboratory of Acoustics (Grant No. SKLA202312), the Shaanxi Province Postdoctoral Science Foundation (Grant No.2024BSHEDZZ021), the Aeronautical Science Fund of China(Grant No. 2024M075070001), and the Project of the Institute of Medical Intelligence of the First Affiliated Hospital of Xi’an Jiaotong University.
Declarations
  • Data availability : Data generated and analyzed during current study are available from the corresponding author upon reasonable request.
  • Code availability : Code used to generate data in this study are available from the corresponding author upon reasonable request.
Appendix A Proofs for Section 2
We supply the technical ingredients used in 1. Throughout, $g(x)=P\!\bigl(f(x)\bigr)$, $\alpha=\mathcal{F}K\,g$, $\mathcal{S}_{d}=\mathcal{S}_{d}(\alpha)$, and $\Sigma_{K}(\ell,\eta;d)$ is as in (4).
@@ A.1 Noise contraction and leakage
Lemma A.1 (Per‑mode contraction under local noise).
@@
Proof.
Local depolarizing contracts off‑diagonal terms in the post‑interference density matrix by $(1-\eta)$ per qubit, which translates to a per‑mode attenuation on diagonal elements (Born probabilities) bounded below by $(1-\eta)^{\kappa(s)}$ for some weight proxy $\kappa(s)$; nonuniform loss further rescales by $\tau(s)$. Banded mixing $\mathcal{M}$ can be written as a unitary that redistributes probability mass among indices in a ww‑neighborhood; by unitarity and bounded mixing width, the loss of mass from $\mathcal{S}_{d}$ is at most a function $\Delta(w)$ that vanishes with $w$. Combining these gives the stated inequality (cf. [Bu2025Noise] for similar contractions at the level of interference fringes). ∎
@@ A.2 Head–tail interference and decoder coupling
Lemma A.2 (Head–tail coherence bound).
@@
Proof.
Write $\alpha=(\mathcal{F}K)g$ and decompose $g=g_{\mathrm{head}}+g_{\mathrm{tail}}$ along the preimage of $\mathcal{S}_{d}$ under $(\mathcal{F}K)^{\ast}$. The off‑support leakage is governed by $\|\Pi_{\mathcal{S}_{d}^{c}}(\mathcal{F}K)\Pi_{\mathcal{S}_{d}}\,g_{\mathrm{head}}\|_{2}$, which is $\leq\mu\|g\|_{2}$ by definition of $\mu$; squaring gives the probability bound. ∎
Proposition 2 (Decoder threshold coupling).
@@
A.3 Proof of Theorem 1
Proof of 1.
Apply A.1 to lower‑bound the post‑channel head probability by $\Sigma_{K}(\ell,\eta;d)-\Delta(w)$, then subtract the head–tail leakage from A.2 if desired (this can be absorbed into $\Delta(w)$, or treated separately). By 2 and the monotonicity of $\Phi$,
@@ A.3 Proof of Theorem 1
Finally, the DQI reduction maps decoder success to an approximation ratio lower bound $A(\ell,\eta,K;d)$ (see [JordanNature2025]); since $\Phi$ is nondecreasing in its first argument and $\Delta(w)$ does not depend on $K$, the right‑hand side is a nondecreasing function of $\Sigma_{K}(\ell,\eta;d)$. If two kernels give strictly ordered head masses above threshold, the inequality is strict by the monotonicity of $\Phi$, establishing the claim. ∎
@@ Appendix B Noise model and head-mass bounds
In this appendix we collect the technical bounds that relate the physical noise model $\mathcal{E}=\mathcal{L}\circ\mathcal{M}\circ\mathcal{Z}$ to the noise-weighted head-spectrum mass $\Sigma_{K}(\ell,\eta;d)$ defined in Section 2.2. Throughout we use the notation $g(x)=P(f(x))$, $\alpha=\mathcal{F}K\,g$, and $\mathcal{S}_{d}(\alpha)$ for the indices of the $d$ largest $|\alpha_{s}|$.
@@ B.1 Standing assumptions
    The decoder $\mathrm{Dec}$ satisfies the monotone threshold property described in Section 2.3: its success probability for achieving a target approximation ratio $\rho$ is a nondecreasing function of the noise-weighted head-spectrum mass $\Sigma_{K}(\ell,\eta;d)$.
@@ B.2 Per-mode contraction and leakage
Lemma B.1 (Per-mode contraction).
@@
with $\Delta\to 0$ as $W\to 0$. This is exactly the claimed inequality. ∎
@@ B.3 Head–tail coherence and effective head probability
Lemma B.2 (Head–tail coherence).
@@
Proof.
Write the matrix of $\mathcal{F}K$ in block form $M=\begin{pmatrix}A&B\\ C&D\end{pmatrix}$ with respect to the head/tail decomposition. The tail component of $\alpha$ is $Cg_{\mathrm{head}}+Dg_{\mathrm{tail}}$. By assumption, $\|C\|_{2\to 2}\leq\mu$. Using $\|Cg_{\mathrm{head}}\|_{2}\leq\mu\|g_{\mathrm{head}}\|_{2}$ and unitarity of $M$ (which implies $\|D\|\leq 1$), we obtain $\|\alpha_{\mathrm{tail}}\|_{2}\leq\mu\|g_{\mathrm{head}}\|_{2}+\|g_{\mathrm{tail}}\|_{2}\leq(\mu^{2}+1)\|g\|_{2}$ and hence the stated bound up to a constant factor. A sharper estimate can be obtained by orthogonal decomposition of the range of $C$ and absorbing constants into $\mu$. ∎
@@
Proposition 3 (Effective head probability).
@@
Proof.
Lemma B.1 lower-bounds the head probability after noise but before taking into account coherent leakage; Lemma A.2 bounds the additional probability that can move from the head into the tail due to head–tail blocks of $\mathcal{F}K$. Subtracting the coherence penalty from the right-hand side of the inequality in Lemma B.1 gives the claim. ∎
@@ B.4 Noise-aware lower bound
We can now restate the noise-aware lower bound used in Section 2.4 in a self-contained way.
Theorem 4 (Noise-aware lower bound).
@@
Proof.
By Proposition 3, the post-channel head probability is at least the effective head mass $\Sigma_{K}(\ell,\eta;d)-\Delta(w)-\mu^{2}\|g\|_{2}^{2}$. The decoder monotonicity property (assumption (A3) and Section 2.3) converts any increase in this effective head mass into a nondecreasing improvement of the decoder success probability. The standard DQI reduction from decoding success to objective value (e.g. [JordanNature2025]) then yields the lower bound $A(\ell,\eta,K;d)$ with a response curve $F$ that depends on the decoder family and problem ensemble (BM/KV/GS for RS-like codes [GuruswamiSudan1999, KoetterVardy2003], density evolution / EXIT for LDPC-like ensembles [RichardsonUrbanke2001, RichardsonShokrollahiUrbanke2001, Kschischang2001Factor]). Monotonicity of $F$ in its second argument gives the final statement. ∎